OpenStudy (lilday12):

Which of the following will form the composite function G(F(x)) shown below? G(F(x)) = 5(x - 7)2 - 4 A. F(x) = 5x - 4 and G(x) = (x - 7)2 B. F(x) = x2 - 4 and G(x) = x - 7 C. F(x) = (x - 7)2 and G(x) = -4 D. F(x) = x - 7 and G(x) = 5x2 - 4

Directrix (directrix):

G(F(x)) = 5(x - 7)2 - 4 I'm thinking that G(x) is a function that squares the values of x, multiples that by 5 and then subtracts 4. G(x) = 5* ( ? )^2 -4 If F(x) = x - 7, then G(F(x)) = G(x-7) = 5* ( x - 7 )^2 -4 @lilday12

Directrix (directrix):

Which option do you think is correct?